OpenStudy (anonymous):

if f(θ)= cot θ = -2, find f(θ+π) HELP ME PLEASE THIS IS REALLY IMPORTANT ITS TRIG PLEASE PLEASE

OpenStudy (anonymous):

The period of \(\cot\theta\) is \(\pi\), so adding \(\pi\) to \(\theta\) gives you the same value. In other words, \(\cot(\theta+\pi)=\cot\theta\).